2x-5 /(x-1)(x-2)(x-5) find out the partial fraction of above function

2x-5/(x-1)(x-2)(x-5)  find out the partial fraction of above function

Grade:12th pass

2 Answers

Arun
4 years ago
Dear student
 
2x-5/(x-1)(x-2)(x-5) = A/(x-1) + B/(x-2) + C/(x-5)
 
A = – ¾
B = 1/3
C =  5/12
 
hence
 
2x-5/(x-1)(x-2)(x-5) = – 3/4(x-1) + 1/3(x-2) + 5/12(x-5)
Vikas TU
4 years ago
Dear student 
Please follow this link 
Good Luck